OpenStudy (anonymous):

Will medal: How do I reflect over the x-axis? Let's say I have an equation: x^2+3x^2+9 and I want to do a vertical reflection on it, what steps would I follow in order to do so?

OpenStudy (anonymous):

f(x)=x^2+3x^2+9 -f(x)=-(x^2+3x^2+9) -f(x)=-x^2-3x^2-9 since reflect over the x-axis means the sign of the y is changed

OpenStudy (anonymous):

for a parabola equation f(x)=x^2+3x^2+9 its vertex is pointing down all positive and if we revolved it to y axis we will have a solid of revolutionj..

OpenStudy (anonymous):

Yes